# The Effects of Multi-Task Learning on ReLU Neural Network Functions

### v3

@@ 1 Introduction
Neural networks are trained to fit data using gradient descent methods and often include a form of regularization called weight decay, which penalizes the $\ell^{2}$ norm of the network weights. We consider weight decay applied only to the input and output weights of the neurons—no regularization is applied to the biases or residual connection. This is a common setting studied frequently in past work Savarese et al. (2019); Ongie et al. (2019); Parhi and Nowak (2021). Intuitively, only the input and output weights—not the biases or residual connection—affect the “regularity” of the neural network function as measured by its second (distributional) derivative, which is why it makes sense to regularize only these parameters. Given a set of training data points $({\bm{x}}_{1},{\bm{y}}_{1}),\dots,({\bm{x}}_{N},{\bm{y}}_{N})\in\mathbb{R}^{d}\times\mathbb{R}^{T}$ and a fixed width $K\geq N^{2}$, we consider the weight decay interpolation problem:  [^1] By an argument similar to the proof of Theorem 5 of Shenouda et al. (2024), as long as $K\geq N^{2}$, problem (3) is well-posed and attains the same minimal objective value (regardless of which $K\geq N^{2}$ is chosen). Therefore, in this work, we always assume that $K$ is some fixed value larger than $N^{2}$.
@@
The optimizations in (2) and (3) are non-convex and in general, they may have multiple global minimizers. As an example, consider the single-task, univariate dataset in Fig. 1. For this dataset, (3) has infinitely many global solutions Savarese et al. (2019); Ergen and Pilanci (2021); Debarre et al. (2022); Hanin (2021). Two of the global minimizers are shown in Fig. 1. In some scenarios, the solution on the right may be preferable to the one on the left, since the interpolation function stays closer to the training data points, and hence is more adversarially robust by most definitions Carlini et al. (2019). Moreover, recent theoretical work shows that this solution has other favorable generalization and robustness properties Joshi et al. (2024). Current training methods, however, might produce any one of the infinite number of solutions, depending on the random initialization of neural network weights as well as other possible sources of randomness in the training process. It is impossible to control this using existing training algorithms, which might explain many problems associated with current neural networks such as their sensitivity to adversarial attacks. In contrast, as we show in this paper, if we train a neural network to interpolate the data in Fig. 1 along with additional interpolation tasks with different labels, then the solution is almost always unique and given by the (potentially preferable) interpolation depicted on the right. This demonstrates that multi-task learning can have a profound effect on the nature of neural network solutions.
@@
    In the univariate setting ($d=1$) we prove that the solution to multi-task learning with different tasks is almost always a unique function, and we give a precise condition for the exceptional cases where solutions are non-unique.
@@
    In the case that the solution to the univariate weight decay problem is unique, it is given by the connect-the-dots interpolant of the training data points: i.e., the optimal solution is a linear spline which performs straight-line interpolation between consecutive data points in all tasks. On the support of the data, this solution agrees with the solution to minimum-norm interpolation in the first-order Sobolev space $H^{1}$, a reproducing kernel Hilbert space (RKHS) which contains all functions with first derivatives in $L^{2}$ De Boor and Lynch (1966). In contrast, solutions to the single-task learning problem are non-unique in general and are given by minimum-norm interpolating functions in the non-Hilbertian Banach space $\operatorname{BV}^{2}$ Parhi and Nowak (2021), which contains all functions with second distributional derivatives [^2] Technically, $\operatorname{BV}^{2}$ contains all functions with second distributional derivatives in ${\mathcal{M}}$, the space of Radon measures with finite total variation. ${\mathcal{M}}$ can be viewed as a “generalization” of $L^{1}$ (see Ch. 7.3, p.223 in Folland (1999)). in $L^{1}$. This shows that the solution to each individual task in a multi-task solution is equivalent to that of a minimum-norm kernel interpolation problem, whereas single-task solutions generally are not.
@@
    We provide empirical evidence and mathematical analysis which indicate that similar conclusions hold in multivariate settings. Specifically, with a large number of tasks, the solutions for each individual task in multi-task learning are approximately minimum-norm solutions in a particular RKHS determined by the optimal neurons. In contrast, learning each task in isolation results in solutions that are minimum-norm with respect to a non-Hilbertian Banach norm over the optimal neurons.
@@ 2 Related Works
Hanin (2021); Stewart et al. (2023) characterized the neural network solutions to (2) in the univariate input/output setting. Boursier and Flammarion (2023) showed that in the univariate input/output case, when weight decay is modified to include the biases of each neuron, the solution is unique. Moreover, under certain assumptions, it is often the sparsest interpolant (i.e., the interpolant with the fewest neurons). Our work differs from these in that we study the multi-task setting, showing that univariate multi-task solutions are almost always unique and equivalent to the connect-the-dots solution, which is generally not the sparsest, and is a minimum-norm solution in a Sobolev RKHS. While characterizing solutions to (2) in the multivariate setting is more challenging, there exist some results for very particular datasets Ergen and Pilanci (2021); Ardeshir et al. (2023); Zeno et al. (2024).
@@
The advantages of multi-task learning has been extensively studied in the machine learning literature Obozinski et al. (2006, 2010); Argyriou et al. (2006, 2008); Caruana (1997). Theoretical properties of multi-task neural networks in particular, have also been studied in Lindsey and Lippl (2023); Collins et al. (2024); Shenouda et al. (2024). The underlying intuition in these past works has been that learning multiple related tasks simultaneously can help select or learn the most useful features for all tasks. Our work differs from this traditional paradigm as we consider multi-task neural networks trained on very general tasks which may be diverse and unrelated.
@@ 3 Univariate Multi-Task Neural Network Solutions
$R(f):=\inf_{\bm{\theta}}\sum_{k=1}^{K}\|{\bm{v}}_{k}\|_{2}\ \quad\text{subject to}\quad\|{\bm{w}}_{k}\|_{2}=1\ \forall k,\quad f=f_{{\bm{\theta}}}$ (4)
@@
denote the slope of the straight line between $(x_{i},y_{it})$ and $(x_{i+1},y_{i+1t})$. The connect-the-dots interpolant of the dataset ${\mathcal{D}}_{t}$ is the function $f_{{\mathcal{D}}_{t}}$ which connects the consecutive points in dataset ${\mathcal{D}}_{t}$ with straight lines (see Fig. 2). Its slopes on $(-\infty,x_{2}]$ and $[x_{N-1},\infty)$ are $s_{1t}$ and $s_{N-1t}$, respectively.
In the following section, we state a simple necessary and sufficient condition under which the connect-the-dots interpolation $f_{{\mathcal{D}}}=(f_{{\mathcal{D}}_{1}},\dots,f_{{\mathcal{D}}_{T}})$ is the unique optimal interpolant of the datasets ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{T}$. We also demonstrate that the set of multi-task datasets which satisfy the necessary condition for non-uniqueness, viewed as a subset of $\mathbb{R}^{N}\times\mathbb{R}^{T\times N}$, has Lebesgue measure zero.
@@
$\min_{\alpha_{1},\dots,\alpha_{N}\in\mathbb{R}}\sum_{i=1}^{N}\sum_{j=1}^{N}\alpha_{i}\alpha_{j}\,k(x_{i},x_{j})\mbox{ \ s.t. \ }\sum_{j=1}^{N}\alpha_{j}\,k(x_{i},x_{j})=y_{it},\quad i=1,\dots,N,$
with the kernel $k(x,x^{\prime})=1-(x-x^{\prime})_{+}+(x-x_{1})_{+}+(x_{1}-x^{\prime})_{+}$ De Boor and Lynch (1966). Therefore, our result shows that the individual outputs of solutions to (3) for $T>1$ tasks almost always coincide on $[x_{1},x_{N}]$ with this kernel solution. In contrast, optimal solutions to the (3) in the case $T=1$ are generally non-unique and may not coincide with the connect-the-dots kernel solution Hanin (2022). We note that for $T=1$, our result is consistent with the characterization of univariate solutions to (3) in Hanin (2022).
@@ 3.1 Characterization and Uniqueness
are both nonzero and aligned [^3] Two vectors ${\bm{u}}_{1}$ and ${\bm{u}}_{2}$ are aligned if ${\bm{u}}_{1}^{\top}{\bm{u}}_{2}=\|{\bm{u}}_{1}\|\|{\bm{u}}_{2}\|$.
@@
When represented by a ReLU neural network each knot in the function corresponds to a neuron. The representational cost (4) is separable across neurons. The contribution of these knots to $R(f)$ is:
@@
Lemma 3.2 states that removing neurons which are located away from the data points and replacing them with a straight line will never increase the representational cost of the network, and it will strictly decrease the representational cost unless ${\bm{a}}-{\bm{b}}$ and ${\bm{b}}-\bm{c}$ are aligned. This result implies that the connect-the-dots interpolant $f_{{\mathcal{D}}}$ is always a solution to (3), since we may take any solution $f$ of (3) and remove all knots from it (resulting in the function $f_{{\mathcal{D}}}$) without increasing its representational cost. If ${\bm{s}}_{i}-{\bm{s}}_{i-1}$ and ${\bm{s}}_{i+1}-{\bm{s}}_{i}$ are aligned for some $i=2,\dots,N-2$, we can view any interpolant on the interval $[x_{i},x_{i+1}]$ as an instance of Fig. 3 with ${\bm{a}}={\bm{s}}_{i-1}$, ${\bm{b}}={\bm{s}}_{i}$, and $\bm{c}={\bm{s}}_{i+1}$. By Lemma 3.2, any neural network with a knot at some point $\tilde{x}\in(x_{i},x_{i+1})$ can have the same representational cost as the connect-the-dots solution on this interval, only if ${\bm{a}}-{\bm{b}}$ and ${\bm{b}}-\bm{c}$ are aligned.
We can also prove by contradiction that optimal solutions are unique on $[x_{i},x_{i+1}]$ as long as ${\bm{s}}_{i}-{\bm{s}}_{i-1}$ and ${\bm{s}}_{i+1}-{\bm{s}}_{i}$ are not aligned. Suppose that there is some other optimal interpolant $f$ which is not the connect-the-dots solution $f_{{\mathcal{D}}}$ on an interval $[x_{i},x_{i+1}]$ for which ${\bm{s}}_{i}-{\bm{s}}_{i-1}$ and ${\bm{s}}_{i+1}-{\bm{s}}_{i}$ are not aligned. Then apply the lemma repeatedly to remove all knots from $f_{\bm{\theta}}$ which are not located at the data points, except for a single remaining knot at some $\tilde{x}$ between consecutive data points. If this knot occurs after $x_{2}$ (the second data point) or before $x_{N-1}$ (second to last data point), the lemma implies automatically (again taking ${\bm{a}}={\bm{s}}_{i-1}$, ${\bm{b}}={\bm{s}}_{i}$, and ${\bm{s}}_{i+1}$) that removing this knot would strictly decrease the representational cost of network, contradicting optimality of $f$. To conclude the proof, it remains only to show that any optimal interpolant of the dataset must agree with the connect-the-dots interpolant $f_{{\mathcal{D}}}$ before $x_{2}$ and after $x_{N-1}$; the details of this argument appear in Section A.1. As our theorem and corollary quantify, real-world regression datasets (which are typically real-valued and often assumed to incorporate some random noise from an absolutely continuous distribution, e.g. Gaussian) are extremely unlikely to satisfy this special alignment condition; hence, our claim that connect-the-dots interpolation is almost always the unique solution to (3).
@@
We provide numerical examples to illustrate the difference in solutions obtained from single task versus multi-task training and validate our theorem. The first row in Fig. 4 shows three randomly initialized univariate neural networks trained to interpolate the five red points with minimum sum of squared weights. While all three of the learned functions have the same representational cost (i.e., all minimize the second-order total variation subject to the interpolation constraint), they each learn different interpolants. This demonstrates that gradient descent does not induce a bias towards a particular solution. The second row shows the function learned for the first output of a multi-task neural network. This network was trained on two tasks. The first task consists of interpolating the five red points while the second consists of interpolating five randomly generated labels sampled from a standard Gaussian distribution. When trained to interpolate with minimum sum of squared weights we see that the connect-the-dots solution is the only solution learned regardless of initialization, verifying Theorem 3.1. This solution simultaneously minimizes the second-order total variation and the norm in the first-order Sobolev RKHS $H^{1}([x_{1},x_{N}])$ associated with the kernel $k(x,x^{\prime})=1-(x-x^{\prime})_{+}+(x-x_{1})_{+}+(x_{1}-x^{\prime})_{+}$ De Boor and Lynch (1966), subject to the interpolation constraints. [^4] Code for reproducing all numerical experiments can be found at https://github.com/joeshenouda/effects-mtl-nns.
Figure 4: Top Row: Three randomly initialized neural networks trained to interpolate the five red points with minimum sum of squared weights. Bottom Row: Three randomly initialized two-output neural networks trained to interpolate a multi-task dataset with minimum sum of squared weights. The labels for the first task are the five red points shown while the labels for the second were randomly sampled from a standard Gaussian distribution.

### v4

@@ 1 Introduction
Neural networks are trained to fit data using gradient descent methods and often include a form of regularization called weight decay, which penalizes the $\ell^{2}$ norm of the network weights. We consider weight decay applied only to the input and output weights of the neurons—no regularization is applied to the biases or residual connection. This is a common setting studied frequently in past work Savarese et al. (2019); Ongie et al. (2019); Parhi and Nowak (2021). Intuitively, only the input and output weights—not the biases or residual connection—affect the “regularity” of the neural network function as measured by its second (distributional) derivative, which is why it makes sense to regularize only these parameters. Given a set of training data points $({\bm{x}}_{1},{\bm{y}}_{1}),\dots,({\bm{x}}_{N},{\bm{y}}_{N})\in\mathbb{R}^{d}\times\mathbb{R}^{T}$ and a fixed width [^1] By an argument similar to the proof of Theorem 5 of Shenouda et al. (2024), as long as $K\geq N^{2}$, problem (3) is well-posed and attains the same minimal objective value (regardless of which $K\geq N^{2}$ is chosen). Therefore, in this work, we always assume that $K$ is some fixed value larger than $N^{2}$. $K\geq N^{2}$, we consider the weight decay interpolation problem:
@@
The optimizations in (2) and (3) are non-convex and in general, they may have multiple global minimizers. As an example, consider the single-task, univariate dataset in Fig. 1. For this dataset, (3) has infinitely many global solutions Savarese et al. (2019); Ergen and Pilanci (2021); Debarre et al. (2022); Hanin (2021). Two of the global minimizers are shown in Fig. 1. In some scenarios, the solution on the right may be preferable to the one on the left, since the interpolation function stays closer to the training data points, and hence is more adversarially robust by most definitions Carlini et al. (2019). Moreover, recent theoretical work shows that this solution has other favorable generalization and robustness properties Joshi et al. (2024). Current training methods, however, might produce any one of the infinite number of solutions, depending on the random initialization of neural network weights as well as other possible sources of randomness in the training process. It is impossible to control this using existing training algorithms, which might explain many problems associated with current neural networks such as their sensitivity to adversarial attacks. In contrast, as we show in this paper, training a network to interpolate the data in Fig. 1 along with additional interpolation tasks with different labels almost always produces a unique solution, given by the (potentially preferable) interpolation depicted on the right. This demonstrates that the solutions to multi-task learning problems can be profoundly different than those of single-task learning problems.
@@
    In the univariate setting ($d=1$) we prove that the solution to multi-task learning problems with different tasks almost always represent a unique function, and we give a precise condition for the exceptional cases where solutions are non-unique.
@@
    When the solution to the univariate weight decay problem is unique, it is given by the connect-the-dots interpolant of the training data points: i.e., the optimal solution is a linear spline which performs straight-line interpolation between consecutive data points in all tasks. On the support of the data, this solution agrees with the minimum-norm interpolant in the first-order Sobolev space $H^{1}$, a reproducing kernel Hilbert space (RKHS) which contains all functions with first derivatives in $L^{2}$ De Boor and Lynch (1966). In contrast, solutions to the single-task learning problem are non-unique in general and are given by minimum-norm interpolating functions in the non-Hilbertian Banach space $\operatorname{BV}^{2}$ Parhi and Nowak (2021), which contains all functions with second distributional derivatives [^2] Technically, $\operatorname{BV}^{2}$ contains all functions with second distributional derivatives in ${\mathcal{M}}$, the space of Radon measures with finite total variation. ${\mathcal{M}}$ can be viewed as a “generalization” of $L^{1}$ (see Ch. 7.3, p.223 in Folland (1999)). in $L^{1}$. This shows that the individual task solutions to a multi-task learning problem are almost always equivalent to those of a minimum-norm kernel interpolation problem, whereas single-task solutions generally are not.
@@
    We provide empirical evidence and mathematical analysis which indicate that similar conclusions hold in multivariate settings. Specifically, the individual task solutions to a multi-task learning problem are approximately minimum-norm solutions in a particular RKHS determined by the optimal neurons. In contrast, learning each task in isolation results in solutions that are minimum-norm with respect to a non-Hilbertian Banach norm over the optimal neurons.
@@ 2 Related Works
Hanin (2021); Stewart et al. (2023) characterized the neural network solutions to (2) in the univariate input/output setting. Boursier and Flammarion (2023) showed that in the univariate input/output case, when weight decay is modified to include the biases of each neuron, the solution is unique. Moreover, under certain assumptions, it is the sparsest interpolant (i.e., the interpolant with the fewest neurons). Our work differs from these in that we study the multi-task setting, showing that univariate-input multi-task solutions are almost always unique and equivalent to the connect-the-dots solution, which is generally not the sparsest, and is a minimum-norm solution in a Sobolev RKHS. While characterizing solutions to (2) in the multivariate setting is more challenging, there exist some results for very particular datasets Ergen and Pilanci (2021); Ardeshir et al. (2023); Zeno et al. (2024).
@@
The advantages of multi-task learning have been extensively studied in the machine learning literature Obozinski et al. (2006, 2010); Argyriou et al. (2006, 2008); Caruana (1997). In particular, the theoretical properties of multi-task neural networks have been studied in Lindsey and Lippl (2023); Collins et al. (2024); Shenouda et al. (2024). The underlying intuition in these past works has been that learning multiple related tasks simultaneously can help select or learn the most useful features for all tasks. Our work differs from this traditional paradigm as we consider multi-task neural networks trained on very general tasks which may be diverse and unrelated.
@@ 3 Univariate Multi-Task Neural Network Solutions
$R(f):=\inf_{\bm{\theta}}\sum_{k=1}^{K}\|{\bm{v}}_{k}\|_{2}\ ,\ \text{subject to }\|{\bm{w}}_{k}\|_{2}=1\ \forall k,\,f=f_{{\bm{\theta}}}$ (4)
@@
denote the slope of the straight line between $(x_{i},y_{it})$ and $(x_{i+1},y_{i+1t})$. The connect-the-dots interpolant of the dataset ${\mathcal{D}}_{t}$ is the function $f_{{\mathcal{D}}_{t}}$ which connects the consecutive points in dataset ${\mathcal{D}}_{t}$ with straight lines (see Fig. 2). Its slopes on $(-\infty,x_{2}]$ and $[x_{N-1},\infty)$ are $s_{1t}$ and $s_{N-1t}$, respectively. In the following section, we state a simple necessary and sufficient condition under which the connect-the-dots interpolation $f_{{\mathcal{D}}}=(f_{{\mathcal{D}}_{1}},\dots,f_{{\mathcal{D}}_{T}})$ is the unique optimal interpolant of the datasets ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{T}$. We also demonstrate that the set of multi-task datasets which satisfy the necessary condition for non-uniqueness, viewed as a subset of $\mathbb{R}^{N}\times\mathbb{R}^{T\times N}$, has Lebesgue measure zero.
@@
$\min_{\alpha_{1},\dots,\alpha_{N}\in\mathbb{R}}\sum_{i=1}^{N}\sum_{j=1}^{N}\alpha_{i}\alpha_{j}k(x_{i},x_{j})\ ,\ \mbox{subject to }\sum_{j=1}^{N}\alpha_{j}k(x_{i},x_{j})=y_{it},\,i=1,\dots,N\ .$ (6)
with the kernel $k(x,x^{\prime})=1-(x-x^{\prime})_{+}+(x-x_{1})_{+}+(x_{1}-x^{\prime})_{+}$ De Boor and Lynch (1966). Therefore, our result shows that the individual outputs of solutions to (3) for $T>1$ tasks almost always coincide on $[x_{1},x_{N}]$ with this kernel solution. In contrast, optimal solutions to the (3) in the case $T=1$ are generally non-unique and may not coincide with the connect-the-dots kernel solution Hanin (2022); Debarre et al. (2022). We note that for $T=1$, our result is consistent with the characterizations of univariate solutions to (3) in Hanin (2022); Debarre et al. (2022).
@@ 3.1 Characterization and Uniqueness
are both nonzero and aligned. [^3] Two vectors ${\bm{u}}_{1}$ and ${\bm{u}}_{2}$ are aligned if ${\bm{u}}_{1}^{\top}{\bm{u}}_{2}=\|{\bm{u}}_{1}\|\|{\bm{u}}_{2}\|$.
@@
When represented by a ReLU neural network, each knot in the function corresponds to a neuron. The representational cost (4) is separable across knots/neurons. The contribution of these knots to $R(f)$ is:
@@
Lemma 3.2 states that removing neurons which are located away from the data points and replacing them with a straight line will never increase the cost of the network, and it will strictly decrease the cost unless ${\bm{a}}-{\bm{b}}$ and ${\bm{b}}-\bm{c}$ are aligned. This result implies that the connect-the-dots interpolant $f_{{\mathcal{D}}}$ is always a solution to (3), since we may take any solution $f$ of (3) and remove all knots from it (resulting in the function $f_{{\mathcal{D}}}$) without increasing its representational cost. If ${\bm{s}}_{i}-{\bm{s}}_{i-1}$ and ${\bm{s}}_{i+1}-{\bm{s}}_{i}$ are aligned for some $i=2,\dots,N-2$, we can view any interpolant on the interval $[x_{i},x_{i+1}]$ as an instance of Fig. 3 with ${\bm{a}}={\bm{s}}_{i-1}$, ${\bm{b}}={\bm{s}}_{i}$, and $\bm{c}={\bm{s}}_{i+1}$. By Lemma 3.2, any CPWL function with a knot at some point $\tilde{x}\in(x_{i},x_{i+1})$ can have the same representational cost as the connect-the-dots solution on this interval, only if ${\bm{a}}-{\bm{b}}$ and ${\bm{b}}-\bm{c}$ are aligned.
We can also prove by contradiction that optimal solutions are unique on $[x_{i},x_{i+1}]$ as long as ${\bm{s}}_{i}-{\bm{s}}_{i-1}$ and ${\bm{s}}_{i+1}-{\bm{s}}_{i}$ are not aligned. Suppose that there is some other optimal interpolant $f$ which is not the connect-the-dots solution $f_{{\mathcal{D}}}$ on an interval $[x_{i},x_{i+1}]$ for which ${\bm{s}}_{i}-{\bm{s}}_{i-1}$ and ${\bm{s}}_{i+1}-{\bm{s}}_{i}$ are not aligned. Then apply the lemma repeatedly to remove all knots from $f_{\bm{\theta}}$ which are not located at the data points, except for a single remaining knot at some $\tilde{x}$ between consecutive data points. If this knot occurs after $x_{2}$ (the second data point) or before $x_{N-1}$ (second to last data point), the lemma implies automatically (again taking ${\bm{a}}={\bm{s}}_{i-1}$, ${\bm{b}}={\bm{s}}_{i}$, and ${\bm{s}}_{i+1}$) that removing this knot would strictly decrease the representational cost of the function, contradicting optimality of $f$. To conclude the proof, it remains only to show that any optimal interpolant of the dataset must agree with the connect-the-dots interpolant $f_{{\mathcal{D}}}$ before $x_{2}$ and after $x_{N-1}$; the details of this argument appear in Section A.1. As our theorem and corollary quantify, real-world regression datasets (which are typically real-valued and often assumed to incorporate some random noise from an absolutely continuous distribution, e.g. Gaussian) are extremely unlikely to satisfy this special alignment condition; hence, our claim that connect-the-dots interpolation is almost always the unique solution to (3).
@@
We provide numerical examples to illustrate the difference in solutions obtained from single task versus multi-task training and validate our theorem. The first row in Fig. 4 shows three randomly initialized univariate neural networks trained to interpolate the five red points with minimum sum of squared weights. While all three of the learned functions have the same representational cost (i.e., all minimize the second-order total variation subject to the interpolation constraint), they each learn different interpolants. This demonstrates that gradient descent does not induce a bias towards a particular solution. The second row shows the function learned for the first output of a multi-task neural network. This network was trained on two tasks. The first task consists of interpolating the five red points while the second consists of interpolating five randomly generated labels sampled from a standard Gaussian distribution. When trained to interpolate with minimum sum of squared weights we see that the connect-the-dots solution is the only solution learned regardless of initialization, verifying Theorem 3.1. This solution simultaneously minimizes the second-order total variation and the norm in the first-order Sobolev RKHS $H^{1}([x_{1},x_{N}])$ associated with the kernel $k(x,x^{\prime})=1-(x-x^{\prime})_{+}+(x-x_{1})_{+}+(x_{1}-x^{\prime})_{+}$ De Boor and Lynch (1966), subject to the interpolation constraints. [^4] The code for all numerical experiments can be found by clicking on the Jupyter logo [Uncaptioned image] in the caption, which links to the notebook that reproduces that figure.
Figure 4: Top Row: Three randomly initialized neural networks trained to interpolate the five red points with minimum sum of squared weights. Bottom Row: Three randomly initialized two-output neural networks trained to interpolate a multi-task dataset with minimum sum of squared weights. The labels for the first task are the five red points shown while the labels for the second were randomly sampled from a standard Gaussian distribution. Refer to caption